Assuming that the atomic nucleus is spherical, calculate the volume of the nucleus given that it's radius is 10^-5 that of the atom with radius 10^-10m.
1
Expert's answer
2021-07-18T01:28:12-0400



The volume of the nucleus will be

"V_

{nucleus}

=

\frac{4}



3

π

⋅

(

10^

{−

5}

)^3

cm^



3"

The volume of the atom will be

"V_

{atom}

=

\frac{4}



3

π

⋅

(

10^{

−

10}
)^3
.

cm



3"

Divide the two volumes to get the fraction of the atom occupied by the nucleus

F

"\frac{V_{nucleus}}{V_{atom}}=\frac{\frac{4}{3}π(10^{-5})^3}
{\frac{4}{3}π(10^{-10})^3}=[\frac{10^{-5}}{10^{-10}}]^3=1.0×10^{15}"


This means that the nucleus occupies "1.0\u00d710^{15}"  of the volume of the atom.